OpenStudy (anonymous):

the nth term of an arithmetic sequence is 4+10n. find the sum of the first 100 terms of the sequence.

OpenStudy (anonymous):

first term should be 14 then \[Sum_{100} = \frac{100(14 + 1004)}{2}\] = 509 * 100 = 50900

OpenStudy (anonymous):

thanks!!